# Duality theory in linear optimization and its extensions — formally verified

### v2

@@ 3.1 Main corollaries
theorem equalityFarkas (A : Matrix I J F) (b : I $\rightarrow$ F) :
[…]
theorem inequalityFarkas [DecidableEq I] (A : Matrix I J F) (b : I $\rightarrow$ F) :
[…]
theorem StandardLP.strongDuality (P : StandardLP I J R) (hP : P.IsFeasible $\vee$ P.dualize.IsFeasible) :
@@ 3.2 Main results
theorem fintypeFarkasBartl {J : Type*} [Fintype J] [LinearOrderedDivisionRing R]
[LinearOrderedAddCommGroup V] [Module R V] [PosSMulMono R V] [AddCommGroup W] [Module R W]
[…]
($\exists$ x : J $\rightarrow$ V, 0 $\leq$ x $\wedge$ $\forall$ w : W, $\sum$ j : J, A w j $\bullet$ x j = b w) $\neq$ ($\exists$ y : W, 0 $\leq$ A y $\wedge$ b y < 0)
[…]
(hAi : $\neg$¬$\exists$ i : I, ($\exists$ j : J, A i j = $\bot$) $\wedge$ ($\exists$ j : J, A i j = $\top$))
(hAj : $\neg$¬$\exists$ j : J, ($\exists$ i : I, A i j = $\bot$) $\wedge$ ($\exists$ i : I, A i j = $\top$))
(hAb : $\neg$¬$\exists$ i : I, ($\exists$ j : J, A i j = $\top$) $\wedge$ b i = $\top$)
(hbA : $\neg$¬$\exists$ i : I, ($\exists$ j : J, A i j = $\bot$) $\wedge$ b i = $\bot$) :
[…]
theorem ValidELP.strongDuality (P : ValidELP I J F) (hP : P.IsFeasible $\vee$ P.dualize.IsFeasible) :
@@ 4 Proving the Farkas-Bartl theorem
Let $n$ be a natural number. Let $R$ be a linearly ordered division ring. Let $W$ be an $R$-module. Let $V$ be a linearly ordered $R$-module whose ordering satisfies monotonicity of scalar multiplication by nonnegative elements on the left. Let $A$ be an $R$-linear map from $W$ to $([\hskip 0.75ptn\hskip 0.75pt]\rightarrow R)$. Let $b$ be an $R$-linear map from $W$ to $V$. Exactly one of the following exists:
@@
  • nonnegative vector family $x:[\hskip 0.75ptn\hskip 0.75pt]\rightarrow V$ such that, for all $w:W$, we have $\sum_{j:[\hskip 0.75ptn\hskip 0.75pt]}\;(A~{}w)_{j}\cdot x_{j}=b~{}w$
  • vector $y:W$ such that $A~{}y\geq 0$ and $b~{}y<0$
@@
The only difference from LABEL:fintypeFarkasBartl is that LABEL:finFarkasBartl uses $[\hskip 0.75ptn\hskip 0.75pt]=\{0,\dots,n\!-\!1\}$ instead of an arbitrary (unordered) finite type $J$.
@@
Proof idea: We first prove that both cannot exist at the same time. Assume we have $x$ and $y$ of said properties. We plug $y$ for $w$ and obtain $\sum_{j:[\hskip 0.75ptn\hskip 0.75pt]}\;(A~{}y)_{j}\cdot x_{j}=b~{}y$. On the left-hand side, we have a sum of nonnegative vectors, which contradicts $b~{}y<0$.
@@
We prove “at least one exists” by induction on $n$. If $n=0$ then $A~{}y\geq 0$ is a tautology. We consider $b$. Either $b$ maps everything to the zero vector, which allows $x$ to be the empty vector family, or some $w$ gets mapped to a nonzero vector, where we choose $y$ to be either $w$ or $(-w)$. Since $V$ is linearly ordered, one of them satisfies $b~{}y<0$. Now we precisely state how the induction step will be.
Lemma (industepFarkasBartl ).
@@
Let $m$ be a natural number. Let $R$ be a linearly ordered division ring. Let $W$ be an $R$-module. Let $V$ be a linearly ordered $R$-module whose ordering satisfies monotonicity of scalar multiplication by nonnegative elements on the left. Assume (induction hypothesis) that for all $R$-linear maps $\bar{A}:W\rightarrow([\hskip 0.75ptm\hskip 0.75pt]\rightarrow R)$ and $\bar{b}:W\rightarrow V$, the formula “$\forall\bar{y}:W\operatorname{,\;}\bar{A}~{}\bar{y}\geq 0\implies\bar{b}~{}\bar{y}\geq 0$” implies existence of a nonnegative vector family $\bar{x}:[\hskip 0.75ptm\hskip 0.75pt]\rightarrow V$ such that, for all $\bar{w}:W$, $\sum_{i:[\hskip 0.75ptm\hskip 0.75pt]}\;(\bar{A}~{}\bar{w})_{i}\cdot\bar{x}_{i}=\bar{b}~{}\bar{w}$. Let $A$ be an $R$-linear map from $W$ to $([\hskip 0.75ptm\!+\!1\hskip 0.75pt]\rightarrow R)$. Let $b$ be an $R$-linear map from $W$ to $V$. Assume that, for all $y:W$, $\;A~{}y\geq 0$ implies $b~{}y\geq 0$. We claim there exists a nonnegative vector family $x:[\hskip 0.75ptm\!+\!1\hskip 0.75pt]\rightarrow V$ such that, for all $w:W$, we have $\sum_{i:[\hskip 0.75ptm+1\hskip 0.75pt]}\;(A~{}w)_{i}\cdot x_{i}=b~{}w$.
@@
Proof idea: Let $A_{<m}$ denote a function that maps $(w:W)$ to $(A~{}w)\big{|}_{[\hskip 0.75ptm\hskip 0.75pt]}$, i.e., $A_{<m}$ is an $R$-linear map from $W$ to $([\hskip 0.75ptm\hskip 0.75pt]\rightarrow R)$ that behaves exactly like $A$ where it is defined. We distinguish two cases. If, for all $y:W$, the inequality $A_{<m}~{}y\geq 0$ implies $b~{}y\geq 0$, then plug $A_{<m}$ for $\bar{A}$, obtain $\bar{x}$, and construct a vector family $x$ such that $x_{m}=0$ and otherwise $x$ copies $\bar{x}$. We easily check that $x$ is nonnegative and that $\sum_{i:[\hskip 0.75ptm+1\hskip 0.75pt]}\;(A~{}w)_{i}\cdot x_{i}=b~{}w$ holds.
@@
In the second case, we have $y^{\prime}$ such that $A_{<m}~{}y^{\prime}\geq 0$ holds but $b~{}y^{\prime}<0$ also holds. We realize that $(A~{}y^{\prime})_{m}<0$. We now declare $y:=((A~{}y^{\prime})_{m})^{-1}\cdot y^{\prime}$ and observe $(A~{}y)_{m}=1$. We establish the following facts (proofs are omitted):
  • for all $w:W$, we have $A~{}(w-((A~{}w)_{m}\cdot y))=0$
  • for all $w:W$, the inequality $A_{<m}~{}(w-((A~{}w)_{m}\cdot y))\geq 0$ implies $b~{}(w-((A~{}w)_{m}\cdot y))\geq 0$
  • for all $w:W$, the inequality $(A_{<m}-(z\mapsto(A~{}z)_{m}\cdot(A_{<m}~{}y)))~{}w\geq 0$ implies $(b-(z\mapsto(A~{}z)_{m}\cdot(b~{}y)))~{}w\geq 0$
@@
We observe that $\bar{A}:=A_{<m}-(z\mapsto(A~{}z)_{m}\cdot(A_{<m}~{}y))$ and $\bar{b}:=b-(z\mapsto(A~{}z)_{m}\cdot(b~{}y))$ are $R$-linear maps. Thanks to the last fact, we can apply induction hypothesis to $\bar{A}$ and $\bar{b}$. We obtain a nonnegative vector family $x^{\prime}$ such that, for all $\bar{w}:W$, $\sum_{i:[\hskip 0.75ptm\hskip 0.75pt]}(\bar{A}~{}\bar{w})_{i}\cdot x^{\prime}_{i}=\bar{b}~{}\bar{w}$. It remains to construct a nonnegative vector family $x:[\hskip 0.75ptm\!+\!1\hskip 0.75pt]\rightarrow V$ such that, for all $w:W$, we have $\sum_{i:[\hskip 0.75ptm+1\hskip 0.75pt]}\;(A~{}w)_{i}\cdot x_{i}=b~{}w$. We choose $x_{m}=b~{}y-\sum_{i:[\hskip 0.75ptm\hskip 0.75pt]}(A_{<m}~{}y)_{i}\cdot x^{\prime}_{i}$ and otherwise $x$ copies $x^{\prime}$. We check that our $x$ has the required properties. ∎
@@ 5 Proving the Extended Farkas theorem
We prove LABEL:inequalityFarkas by applying LABEL:equalityFarkas to the matrix $(1~{}|~{}A)$ where $1$ is the identity matrix of type $(I\times I)\rightarrow F$.
Theorem (inequalityFarkas_neg ).
@@
Theorem (extendedFarkas, restated ).
@@ 6 Proving the Extended strong LP duality
Lemma (unbounded_of_reaches_le ).
@@
Lemma (unbounded_of_feasible_of_neg ).
@@
Lemma (unbounded_of_feasible_of_infeasible ).
@@
Lemma (infeasible_of_unbounded ).
@@
Lemma (dualize_dualize ).
@@
Proof idea: $-(-A^{T})^{T}=A$
Lemma (strongDuality_aux ).
@@
In the first case, we obtain a nonnegative vectors $x$ and $y$ such that $A\cdot x\leq b$, $-A^{T}\cdot y\leq c$, and $c\cdot x+b\cdot y\leq 0$. We immediately see that $P$ reaches $c\cdot x$, that the dual of $P$ reaches $b\cdot y$, and that the desired inequality holds.
In the second case, we obtain nonnegative vectors $y$ and $x$ and a nonnegative scalar $z$ such that $-\!A^{T}\cdot y+z\cdot(-c)\leq 0$, $A\cdot x+z\cdot(-b)\leq 0$, and $b\cdot y+c\cdot x<0$. If $z>0$, we finish the proof using $p:=z^{-1}\cdot c\cdot x$ and $q:=z^{-1}\cdot b\cdot y$ (and we do not care that this case cannot happen in reality). It remains to prove that $z=0$ cannot happen. At least one of $b\cdot y$ or $c\cdot x$ must be strictly negative. If $c\cdot x<0$, we apply LABEL:unbounded-of-feasible-of-neg to $P$, which (using LABEL:infeasible-of-unbounded) contradicts that the dual of $P$ is feasible. If $b\cdot y<0$, we apply LABEL:unbounded-of-feasible-of-neg to the dual of $P$, which (using LABEL:infeasible-of-unbounded and LABEL:dualize-dualize) contradicts that $P$ is feasible.
Lemma (strongDuality_of_both_feasible ).
Let $P$ be a valid linear program such that $P$ is feasible and the dual of $P$ is also feasible. There is a finite value $r$ such that $P$ reaches $-r$ and the dual of $P$ reaches $r$.
@@
Lemma (optimum_unique ).
@@
Proof idea: We prove $r\leq s$ by applying “$P$ is bounded by $r$” to “$s$ is reached by $P$”. We prove $s\leq r$ by applying “$P$ is bounded by $s$” to “$r$ is reached by $P$”.
Lemma (optimum_eq_of_reaches_bounded ).
@@
Lemma (strongDuality_of_prim_feasible ).
@@
Proof idea: If the dual of $P$ is feasible as well, use LABEL:strongDuality-of-both-feasible to obtain $r$ such that $P$ reaches $-r$ and the dual of $P$ reaches $r$. Using LABEL:optimum-eq-of-reaches-bounded together with LABEL:weakDuality, conclude that the optimum of $P$ is $-r$. Using LABEL:optimum-eq-of-reaches-bounded together with LABEL:weakDuality, conclude that the optimum of the dual of $P$ is $r$. Observe that $-r$ and $r$ are opposites.
@@
Theorem (optimum_neq_none ).
@@
Lemma (strongDuality_of_dual_feasible ).
@@
Proof idea: Apply LABEL:strongDuality-of-prim-feas to the dual of $P$ and use LABEL:dualize-dualize.
Theorem (strongDuality, restated ).

### v3

@@ 3.1 Main corollaries
theorem equalityFarkas {I J F : Type*} [Fintype I] [Fintype J]
[LinearOrderedField F] (A : Matrix I J F) (b : I F) :
(∃ x : J F, 0 x A *ᵥ x = b) (∃ y : I F, 0 A *ᵥ y b ⬝ᵥ y < 0)
theorem inequalityFarkas {I J F : Type*} [DecidableEq I] [Fintype I] [Fintype J]
[LinearOrderedField F] (A : Matrix I J F) (b : I F) :
(∃ x : J F, 0 x A *ᵥ x b) (∃ y : I F, 0 y 0 A *ᵥ y b ⬝ᵥ y < 0)
theorem StandardLP.strongDuality {I J R : Type*} [DecidableEq I] [DecidableEq J]
[Fintype I] [Fintype J] [LinearOrderedField R]
(P : StandardLP I J R) (hP : P.IsFeasible P.dualize.IsFeasible) :
@@ 3.2 Main results
theorem fintypeFarkasBartl {J R V W : Type*} [Fintype J] [LinearOrderedDivisionRing R]
[LinearOrderedAddCommGroup V] [Module R V] [PosSMulMono R V]
[AddCommGroup W] [Module R W]
(A : W →ₗ[R] J R) (b : W →ₗ[R] V) :
(∃ x : J V, 0 x w : W, j : J, A w j x j = b w) (∃ y : W, 0 A y b y < 0)
Note that l[R] encodes a linear map as a special case of a semi-linear map [Semilinear].
theorem extendedFarkas {I J F : Type*} [DecidableEq I]
[Fintype I] [Fintype J] [LinearOrderedField F]
(A : Matrix I J F∞) (b : I F∞)
(hAi : ¬∃ i : I, (∃ j : J, A i j = ⊥) (∃ j : J, A i j = ⊤))
(hAj : ¬∃ j : J, (∃ i : I, A i j = ⊥) (∃ i : I, A i j = ⊤))
(hAb : ¬∃ i : I, (∃ j : J, A i j = ⊤) b i = ⊤)
(hbA : ¬∃ i : I, (∃ j : J, A i j = ⊥) b i = ⊥) :
(∃ x : J F≥0, A ₘ* x b) (∃ y : I F≥0, -A ₘ* y 0 b ᵥ⬝ y < 0)
theorem ValidELP.strongDuality {I J F : Type*} [DecidableEq I] [DecidableEq J]
[Fintype I] [Fintype J] [LinearOrderedField F]
(P : ValidELP I J F) (hP : P.IsFeasible P.dualize.IsFeasible) :
@@ 4 Proving the Farkas-Bartl theorem
Let $n$ be a natural number. Let $R$ be a linearly ordered division ring. Let $W$ be an $R$-module. Let $V$ be a linearly ordered $R$-module whose ordering satisfies monotonicity of scalar multiplication by nonnegative elements on the left. Let $A$ be an $R$-linear map from $W$ to $([\thinspace n\thinspace]\rightarrow R)$. Let $b$ be an $R$-linear map from $W$ to $V$. Exactly one of the following exists:
@@
  • nonnegative vector family $x:[\thinspace n\thinspace]\rightarrow V$ such that, for all $w:W$, we have $\sum_{j:[\thinspace n\thinspace]}\;(A\penalty 10000\ w)_{j}\cdot x_{j}=b\penalty 10000\ w$
  • vector $y:W$ such that $A\penalty 10000\ y\geq 0$ and $b\penalty 10000\ y<0$
@@
The only difference from LABEL:fintypeFarkasBartl is that LABEL:finFarkasBartl uses $\thinspace n\thinspace]=\{0,\dots,n\!-\!1\}$ instead of an arbitrary (unordered) finite type $J$.
@@
Proof idea: We first prove that both cannot exist at the same time. Assume we have $x$ and $y$ of said properties. We plug $y$ for $w$ and obtain $\sum_{j:[\thinspace n\thinspace]}\;(A\penalty 10000\ y)_{j}\cdot x_{j}=b\penalty 10000\ y$. On the left-hand side, we have a sum of nonnegative vectors, which contradicts $b\penalty 10000\ y<0$.
@@
We prove “at least one exists” by induction on $n$. If $n=0$ then $A\penalty 10000\ y\geq 0$ is a tautology. We consider $b$. Either $b$ maps everything to the zero vector, which allows $x$ to be the empty vector family, or some $w$ gets mapped to a nonzero vector, where we choose $y$ to be either $w$ or (w)(-w). Since $V$ is linearly ordered, one of them satisfies $b\penalty 10000\ y<0$. Now we precisely state how the induction step will be.
Lemma 4.2 (industepFarkasBartl ).
@@
Let $m$ be a natural number. Let $R$ be a linearly ordered division ring. Let $W$ be an $R$-module. Let $V$ be a linearly ordered $R$-module whose ordering satisfies monotonicity of scalar multiplication by nonnegative elements on the left. Assume (induction hypothesis) that for all $R$-linear maps $\bar{A}:W\rightarrow([\thinspace m\thinspace]\rightarrow R)$ and $\bar{b}:W\rightarrow V$, the formula $\forall\bar{y}:W\operatorname{,\;}\bar{A}\penalty 10000\ \bar{y}\geq 0\implies\bar{b}\penalty 10000\ \bar{y}\geq 0$” implies existence of a nonnegative vector family $\bar{x}:[\thinspace m\thinspace]\rightarrow V$ such that, for all $\bar{w}:W$, $\sum_{i:[\thinspace m\thinspace]}\;(\bar{A}\penalty 10000\ \bar{w})_{i}\cdot\bar{x}_{i}=\bar{b}\penalty 10000\ \bar{w}$. Let $A$ be an $R$-linear map from $W$ to $([\thinspace m\!+\!1\thinspace]\rightarrow R)$. Let $b$ be an $R$-linear map from $W$ to $V$. Assume that, for all $y:W$, $A\penalty 10000\ y\geq 0$ implies $b\penalty 10000\ y\geq 0$. We claim there exists a nonnegative vector family $x:[\thinspace m\!+\!1\thinspace]\rightarrow V$ such that, for all $w:W$, we have $\sum_{i:[\thinspace m+1\thinspace]}\;(A\penalty 10000\ w)_{i}\cdot x_{i}=b\penalty 10000\ w$.
@@
Proof idea: Let $A_{<m}$ denote a function that maps $(w:W)$ to $(A\penalty 10000\ w)\big|_{[\thinspace m\thinspace]}$, i.e., $A_{<m}$ is an $R$-linear map from $W$ to $([\thinspace m\thinspace]\rightarrow R)$ that behaves exactly like $A$ where it is defined. We distinguish two cases. If, for all $y:W$, the inequality $A_{<m}\penalty 10000\ y\geq 0$ implies $b\penalty 10000\ y\geq 0$, then plug $A_{<m}$ for $\bar{A}$, obtain $\bar{x}$, and construct a vector family $x$ such that $x_{m}=0$ and otherwise $x$ copies $\bar{x}$. We easily check that $x$ is nonnegative and that $\sum_{i:[\thinspace m+1\thinspace]}\;(A\penalty 10000\ w)_{i}\cdot x_{i}=b\penalty 10000\ w$ holds.
@@
In the second case, we have $y^{\prime}$ such that $A_{<m}\penalty 10000\ y^{\prime}\geq 0$ holds but $b\penalty 10000\ y^{\prime}<0$ also holds. We realize that $(A\penalty 10000\ y^{\prime})_{m}<0$. We now declare $y:=((A\penalty 10000\ y^{\prime})_{m})^{-1}\cdot y^{\prime}$ and observe $(A\penalty 10000\ y)_{m}=1$. We establish the following facts (proofs are omitted):
  • for all $w:W$, we have $A\penalty 10000\ (w-((A\penalty 10000\ w)_{m}\cdot y))=0$
  • for all $w:W$, the inequality $A_{<m}\penalty 10000\ (w-((A\penalty 10000\ w)_{m}\cdot y))\geq 0$ implies $b\penalty 10000\ (w-((A\penalty 10000\ w)_{m}\cdot y))\geq 0$
  • for all $w:W$, the inequality $(A_{<m}-(z\mapsto(A\penalty 10000\ z)_{m}\cdot(A_{<m}\penalty 10000\ y)))\penalty 10000\ w\geq 0$ implies
    $(b-(z\mapsto(A\penalty 10000\ z)_{m}\cdot(b\penalty 10000\ y)))\penalty 10000\ w\geq 0$
@@
We observe that $\bar{A}:=A_{<m}-(z\mapsto(A\penalty 10000\ z)_{m}\cdot(A_{<m}\penalty 10000\ y))$ and $\bar{b}:=b-(z\mapsto(A\penalty 10000\ z)_{m}\cdot(b\penalty 10000\ y))$ are $R$-linear maps. Thanks to the last fact, we can apply induction hypothesis to $\bar{A}$ and $\bar{b}$. We obtain a nonnegative vector family $x^{\prime}$ such that, for all $\bar{w}:W$, $\sum_{i:[\thinspace m\thinspace]}(\bar{A}\penalty 10000\ \bar{w})_{i}\cdot x^{\prime}_{i}=\bar{b}\penalty 10000\ \bar{w}$. It remains to construct a nonnegative vector family $x:[\thinspace m\!+\!1\thinspace]\rightarrow V$ such that, for all $w:W$, we have $\sum_{i:[\thinspace m+1\thinspace]}\;(A\penalty 10000\ w)_{i}\cdot x_{i}=b\penalty 10000\ w$. We choose $x_{m}=b\penalty 10000\ y-\sum_{i:[\thinspace m\thinspace]}(A_{<m}\penalty 10000\ y)_{i}\cdot x^{\prime}_{i}$ and otherwise $x$ copies $x^{\prime}$. We check that our $x$ has the required properties. ∎
@@ 5 Proving the Extended Farkas theorem
We prove LABEL:inequalityFarkas by applying LABEL:equalityFarkas to the matrix $(1\penalty 10000\ |\penalty 10000\ A)$ where $1$ is the identity matrix of type $(I\times I)\rightarrow F$.
Theorem 5.1 (inequalityFarkas_neg ).
@@
Theorem 5.2 (extendedFarkas, restated ).
@@ 6 Proving the Extended strong LP duality
Lemma 6.4 (unbounded_of_reaches_le ).
@@
Lemma 6.5 (unbounded_of_feasible_of_neg ).
@@
Lemma 6.6 (unbounded_of_feasible_of_infeasible ).
@@
Lemma 6.7 (infeasible_of_unbounded ).
@@
Lemma 6.8 (dualize_dualize ).
@@
Proof idea: $(-A^{T})^{T}=A$
Lemma 6.9 (strongDuality_aux ).
@@
In the first case, we obtain a nonnegative vectors $x$ and $y$ such that $A\cdot x\leq b$, $A^{T}\cdot y\leq c$, and $c\cdot x+b\cdot y\leq 0$. We immediately see that $P$ reaches $c\cdot x$, that the dual of $P$ reaches $b\cdot y$, and that the desired inequality holds.
In the second case, we obtain nonnegative vectors $y$ and $x$ and a nonnegative scalar $z$ such that $A^{T}\cdot y+z\cdot(-c)\leq 0$, $A\cdot x+z\cdot(-b)\leq 0$, and $b\cdot y+c\cdot x<0$. If $z>0$, we finish the proof using $p:=z^{-1}\cdot c\cdot x$ and $q:=z^{-1}\cdot b\cdot y$ (and we do not care that this case cannot happen in reality). It remains to prove that $z=0$ cannot happen. At least one of $b\cdot y$ or $c\cdot x$ must be strictly negative. If $c\cdot x<0$, we apply LABEL:unbounded-of-feasible-of-neg to $P$, which (using LABEL:infeasible-of-unbounded) contradicts that the dual of $P$ is feasible. If $b\cdot y<0$, we apply LABEL:unbounded-of-feasible-of-neg to the dual of $P$, which (using LABEL:infeasible-of-unbounded and LABEL:dualize-dualize) contradicts that $P$ is feasible.
Lemma 6.10 (strongDuality_of_both_feasible ).
Let $P$ be a valid linear program such that $P$ is feasible and the dual of $P$ is also feasible. There is a finite value $r$ such that $P$ reaches r-r and the dual of $P$ reaches $r$.
@@
Lemma 6.11 (optimum_unique ).
@@
Proof idea: We prove $r\leq s$ by applying “$P$ is bounded by rr” to “$s$ is reached by PP”. We prove $s\leq r$ by applying “$P$ is bounded by ss” to “$r$ is reached by PP”.
Lemma 6.12 (optimum_eq_of_reaches_bounded ).
@@
Lemma 6.13 (strongDuality_of_prim_feasible ).
@@
Proof idea: If the dual of $P$ is feasible as well, use LABEL:strongDuality-of-both-feasible to obtain $r$ such that $P$ reaches r-r and the dual of $P$ reaches $r$. Using LABEL:weakDuality together with LABEL:optimum-eq-of-reaches-bounded, conclude that the optimum of $P$ is r-r. Using LABEL:weakDuality together with LABEL:optimum-eq-of-reaches-bounded, conclude that the optimum of the dual of $P$ is $r$. Observe that r-r and $r$ are opposites.
@@
Theorem 6.14 (optimum_neq_none ).
@@
Lemma 6.15 (strongDuality_of_dual_feasible ).
@@
Proof idea: Apply LABEL:strongDuality-of-prim-feas to the dual of $P$ and finish the proof using LABEL:dualize-dualize.
Theorem 6.16 (strongDuality, restated ).